# On Memorization of Large Language Models in Logical Reasoning

### v1

@@ Abstract
Large language models (LLMs) achieve good performance on challenging reasoning benchmarks, yet could also make basic reasoning mistakes. This contrasting behavior is puzzling when it comes to understanding the mechanisms behind LLMs’ reasoning capabilities. One hypothesis is that the increasingly high and nearly saturated performance on common reasoning benchmarks could be due to the memorization of similar problems. In this paper, we systematically investigate this hypothesis with a quantitative measurement of memorization in reasoning tasks, using a dynamically generated logical reasoning benchmark based on Knights and Knaves (K&K) puzzles. We found that LLMs could interpolate the training puzzles (achieving near-perfect accuracy) after fine-tuning, yet fail when those puzzles are slightly perturbed, suggesting that the models heavily rely on memorization to solve those training puzzles. On the other hand, we show that while fine-tuning leads to heavy memorization, it also consistently improves generalization performance. In-depth analyses with perturbation tests, cross difficulty-level transferability, probing model internals, and fine-tuning with wrong answers suggest that the LLMs learn to reason on K&K puzzles despite training data memorization. This phenomenon indicates that LLMs exhibit a complex interplay between memorization and genuine reasoning abilities. Finally, our analysis based on a per-sample memorization score sheds light on how LLMs switch between reasoning and memorization when solving logical puzzles. Our code and data are available at https://memkklogic.github.io/.
@@ 1 Introduction
Modern Large Language Models (LLMs) show impressive reasoning capabilities that allow them to solve a wide range of challenging problems including commonsense reasoning and mathematical reasoning. In the meantime, LLMs also make mistakes on some of the most basic problems, such as comparing which number is bigger—13.11 or 13.8 (Lin, 2024), and counting the number of sisters that Alice’s brother has (Nezhurina et al., 2024). This contrast is puzzling when it comes to understanding how exactly LLMs solve reasoning tasks. This question is important both scientifically and practically: understanding how LLMs reason could shed light on their learning and generalization behaviors. It is also crucial for real-world applications where robust reasoning is required due to safety and trustworthiness concerns (Wang et al., 2023, Wallace et al., 2024, Lee et al., 2024, Wei et al., 2024a).
@@
One hypothesis is that LLMs could be relying on memorization when solving those reasoning tasks, especially when measured by popular benchmarks that could be accidentally leaked into various massive internet-crawled pre-training datasets. Previous work (Tirumala et al., 2022, Carlini et al., 2023) show that LLMs could indeed memorize the training data, which may lead to potential privacy (Carlini et al., 2021) or copyright (Karamolegkou et al., 2023, Wei et al., 2024b) concerns. Additional evidences of potential memorization come from extensive studies on data contamination in LLMs (Magar & Schwartz, 2022, Balloccu et al., 2024, Shi et al., 2024, Xu et al., 2024, Oren et al., 2024). To mitigate the issue of benchmark saturation potentially due to memorization, some papers focus on designing dynamic benchmarks (Roberts et al., 2023, Zhu et al., 2024, Srivastava et al., 2024, Jain et al., 2024, Wu et al., 2024a) or alternative evaluation protocols (Zeng et al., 2023, Zhang et al., 2024, Xu et al., 2024, Srivastava et al., 2024).
@@
In this paper, we take a direct approach to quantify the memorization behaviors of LLMs in reasoning tasks within a controlled setting. Specifically, we seek to understand: (i) whether LLMs rely on memorization to solve reasoning tasks, and (ii) whether memorization is only detrimental to learning to reason. Both questions are inspired by human behavior. For instance, when a student works hard on the preparation material for an exam, the preparation could help them get familiarized with the problems, and their ability to solve new problems could usually improve with enough exercises. However, without genuinely understanding the principles, they might fail when the same problem is slightly changed despite doing well on prepared problems. Our metric of memorization, illustrated in Fig. 1, is based on this intuition. We note that a similar perturbation (mostly at language-level) idea has been used in previous work, especially in detecting contamination (Golchin & Surdeanu, 2023, Yang et al., 2023, Xu et al., 2024). However, given our focus on understanding memorization in logical reasoning tasks, we further consider problem-level perturbation that slightly changes the mathematical structure of a puzzle, in addition to language-level perturbations. To facilitate our study, we propose a new logical reasoning benchmark that supports automatic problem-level perturbation. With this tool, we evaluate 11 off-the-shelf models, and fine-tuned Llama3-8B and GPT4o-mini to quantify memorization in reasoning tasks, and reveal interesting interplay between memorization and reasoning: while models indeed tend to memorize many logical puzzles, they develop genuine reasoning capabilities during fine-tuning (even directly on question-answer pairs without reasoning steps), and the reasoning performance improves as the memorization level increases.
Figure 1: Illustration of the definition of Local Inconsistency based Memorization Score, $\mathsf{LiMem}$.
In the following, we provide an outline of the paper and summarize our key contributions:
@@
  • To facilitate the measurement, we propose a new logical reasoning benchmark based on the Knights and Knaves (K&K, Smullyan, 1978, Johnson-Laird & Byrne, 1990) puzzles, that can generate new puzzles at different difficulty levels, locally perturb existing puzzles, and automatically synthesize detailed reasoning steps to solve a given puzzle (§ 2.2).
@@
  • We design various in-depth analyses (§ 4.1$\sim$ 4.3) to verify that the models developed genuine reasoning capabilities after fine-tuning even with only question-answer pairs and wrong answers, via local perturbation tests, cross difficulty-level transferability, and model internal probing.
@@
2 How to Measure Memorization in Reasoning Tasks
@@ 2.1 Memorization Metrics for Reasoning Tasks
Memorization of LLMs has been studied in various contexts such as privacy (Carlini et al., 2023), copyright  (Carlini et al., 2021, Karamolegkou et al., 2023, Wei et al., 2024b, He et al., 2024), and solving knowledge intensive tasks (Hartmann et al., 2023). In this paper, we are specifically interested in measuring the level of memorization when solving reasoning tasks, by borrowing intuition from human behavior. For example, when preparing for an exam, a student may not be able to fully digest the underlying principles due to various reasons or constraints. But when (luckily) facing the same problem the student had prepared for, they would still be able to solve it. A key characteristic of this type of memorization is: (A) high accuracy on observed problems and (B) low accuracy when the problem is slightly changed. Based on this intuition, for a dataset $\mathcal{D}$ of reasoning puzzles, we combine the following two quantities to measure memorization:
@@
    For (A), we measure the accuracy of a target model $f$ on $\mathcal{D}$, denoted as $\mathsf{Acc}(f;\mathcal{D})$. We are especially interested in measuring on the set of observed puzzles, i.e., the training set, $\mathsf{Acc}(f;\mathsf{Tr})$. We say $f$ interpolates (Belkin et al., 2018, Muthukumar et al., 2020, Belkin, 2021, Bartlett et al., 2021) the training puzzles if $\mathsf{Acc}(f;\mathsf{Tr})\approx 100\%$.
@@
We combine the two factors to define a Local Inconsistency-based Memorization Score:
$\mathsf{LiMem}(f;\mathcal{D})=\mathsf{Acc}(f;\mathcal{D})\cdot(1-\mathsf{CR}(f;\mathcal{D})).$ (1)
@@
When there is no ambiguity, we simply call it the memorization score. $\mathsf{LiMem}(f;\mathcal{D})\in[0,1]$ and a larger score provides stronger evidence of memorization. Specifically, a high $\mathsf{LiMem}(f;\mathsf{Tr})$ matches the characteristic behavior of human memorizing observed puzzles, and in this case we say $f$ memorized the training puzzles. Furthermore, we also measure $\mathsf{LiMem}(f;\mathsf{Tst})$ on test examples, to study if the generalization accuracy is due to reasoning or memorization. Note the $\mathsf{Acc}(f;\mathcal{D})$ factor is simple but necessary, as there are three types of behaviors: (i) solving by memorization, (ii) solving by reasoning, (iii) not solving (e.g., random guessing). A high $\mathsf{LiMem}(f;\mathcal{D})$ indicates (i), but a low $\mathsf{LiMem}(f;\mathcal{D})$ would only indicate (ii) if we separately check that $\mathsf{Acc}(f;\mathcal{D})$ is high.
@@
To effectively measure the memorization score $\mathsf{LiMem}(f;\mathcal{D})$, we need a principled way to (1) perform a local perturbation that changes the puzzle while maintaining its difficulty level; (2) compute the new correct answer after perturbation. Towards this goal, we design and implement a functional dataset based on the Knights and Knaves puzzles (Smullyan, 1978, Johnson-Laird & Byrne, 1990).
2.2 Knights and Knaves Logical Reasoning Benchmark
Figure 2: K&K data generation framework employs abstract and natural language modules to generate question answer pair and synthetic CoTs for each K&K sample, based on the problem specification: number of persons ($N$), tree width ($W$), and depth ($D$). Perturbers in these modules can alter the math structure and language description, respectively.
Knights and Knaves (K&K) is a type of logical puzzle where some characters can only answer questions truthfully, and others only falsely. The goal is to infer each character’s truthfulness. An example of the puzzle and its corresponding answer is provided in Fig. 2.
Based on the K&K puzzle, we design a dynamic benchmark that supports generating new puzzles and perturbing existing puzzles. Our library automatically solves the K&K puzzles and generates detailed reasoning steps and solutions for evaluation and training. Moreover, to support measuring memorization, we also provide a procedure to perturb a given puzzle and recompute the new solution after the perturbation. Specifically, our benchmark consists of two modules:
The Abstract Module has four components that can generate and manipulate K&K puzzles in an abstract form (see § A.1). (1) The Generator creates a random K&K puzzle given a specification $(N,D,W)$ that determines the difficulty level. Specifically, it generates a puzzle with $N$ people, and for each person, a statement that consists of a random tree of maximum width $W$ and depth $D$, where each node is sampled uniformly at random from the following candidate sets. The leaf node can be a claim that a specific person is lying (i.e., knave) or truth-telling (i.e., knight), and the branching node can be and, or, not, imply, and equivalence. (2) The Solver finds the solution to a given puzzle. In our study, we ignore puzzles with no or more than one solution. So we implement the solver by converting the puzzle to a Boolean satisfiability problem and enumerate all Boolean assignments so that we can easily obtain a list of all valid solutions to filter out unwanted puzzles. (3) The Reasoner generates a step-by-step reasoning procedure that leads to the solution. We design it to mimic the reasoning steps used by humans and some LLMs: instead of enumerating all Boolean assignments, it examines each person sequentially, makes an assumption (knight/knave) and checks if it leads to a contradiction with statements from people with specific identities assumed. If not, it continues to examine the next person; otherwise it will try an alternative assumption or backtrack to a previously examined person (details in § A.3). (4) The Perturber, given a puzzle, generates a locally perturbed version that is (superficially) similar to the original puzzle, and solvable with the same underlying principle and at a similar difficulty level. The Perturber replaces either an entire statement or a leaf node in a statement with a newly sampled one. The process is rerun until the perturbed puzzle has a unique solution different from the original puzzle, or until a maximum number of attempts is reached. This rarely happens for $N\geq 3$ people puzzles. When it happens we skip the puzzle in our perturbation analysis. See Tab. 1 for concrete examples.
The Natural Language Module has three components that operate in natural language space. (1) The NL-Generator takes an abstract K&K puzzle and formats it in natural language. It is template-based, with randomly sampled person names and a random template for making claims, and it uses a few heuristics to convert tree-structured logical statements to natural language. (2) The NL-Reasoner converts the reasoning steps computed by the abstract Reasoner to the natural language format in a similar manner. See Fig. 15 for dataset length distributions. (3) The NL-Perturber generates perturbed puzzles by keeping the abstract puzzle intact and manipulating the language-level descriptions as follows (See § A.7): (i) replace character names with uncommon names; (ii) replace knight/knave with other similar pairs of role names, e.g., saint/sinner; (iii) reorder the statements from the characters; (iv) flip the role name from knight/knave to knave/knight. Note that the flipped role perturbation is somewhat adversarial as it goes against the common intuition that a good character would tell the truth and a bad one would lie, so we include it mostly for reference purposes.
3 Quantifying LLM Memorization in Reasoning Tasks
We measure memorization for off-the-shelf models (§ 3.1) and fine-tuned models (§ 3.2).
3.1 Off-the-shelf Models
Evaluation setup. To generate our K&K benchmark (§ 2.2), we use the max tree width $W=2$ and depth $D=2$, and create 100 test puzzles for each $N$-people task ($N\in\{2,3,\ldots,8\}$). Then, we generate perturbed versions for each puzzle under 6 perturbation types introduced in § 2.2: {perturbed statement, perturbed leaf node, random role-pair name, uncommon person name, reordered statement, flipped role}. We utilize 0-shot direct prompting with task-specific instructions for open-ended question-answering (details in § B.2). Note that even under direct prompting, capable LLMs can generate Chain of Thought (Wei et al., 2022, CoT). Our evaluation mainly considers the 00-shot setting, excluding potential biases introduced by in-context examples (Zhao et al., 2021). We defer the results under CoT prompting (i.e., explicitly adding a CoT trigger “Let’s think step by step”), and 1-shot prompting to § C. To evaluate the accuracy of the model’s output, we use keyword matching for full-puzzle correctness. A response is considered correct if every person’s true identity is included in the conclusion part of the model’s response.
Off-the-shelf models do not perform well on K&K tasks. We evaluate 11 models that are shown to perform competitively on common reasoning benchmarks [^1] We are not evaluating the OpenAI o1 model because API access is limited to only the highest-tier users.. As shown in Fig. 3, K&K benchmark poses a challenging logical reasoning task for all the models. Even for the easiest puzzles involving only $2$ persons, the best models still achieve $\leq 70\%$ accuracy. And the performance drops significantly as the complexity increases (the best accuracy is only $11\%$ for $8$-people puzzles).
Figure 3: Test accuracy $\mathsf{Acc}(f;\mathsf{Tst})$ of off-the-shelf models under 0-shot direct prompting drops with increasing puzzle complexity (left). $\mathsf{LiMem}(f;\mathsf{Tst})$ on test examples under statement perturbation (middle) and leaf perturbation (right) is large for specific models, indicating signs of memorization in solving these puzzles.
[…]
3.2 Fine-tuned Models
Fine-tuning setup. We take Llama3-8B and GPT4o-mini and run supervised fine-tuning (SFT) on a set of K&K training puzzles disjoint from the test set. We consider two fine-tuning paradigms: (1) Fine-tuning on detailed CoT steps (CoT FT): during SFT, the model observes the concatenation of the question, synthetic CoT steps, and the answer for each puzzle; the loss is computed on the CoT steps and the answer part. (2) Fine-tuning on the answers (Direct FT) where the model observes the question-answer pair for each puzzle, and the loss is only computed on the answer part. Examples of CoT FT/Direct FT training instances are provided in § B.2.2. We fine-tune the models for each $N$-people task separately, with $n_{\text{train}}=1,000$ for $3\leq N\leq 8$, and $n_{\text{train}}=200$ for $2$-people task due to limited number of combinations. We fine-tune Llama3-8B for $100$ epochs and GPT4o-mini for $5$ epochs (due to budget constraints) via the OpenAI fine-tune API (details in § B.2). During the evaluation, we follow the same prompting paradigm as FT paradigm, i.e., direct/CoT prompting for direct/CoT-FTed model, which is shown effective in § C.2.
LLMs interpolate K&K training puzzles. In Fig. 5, we present the training accuracy of models trained on each task on the $x$-axis (each dot represents a training epoch). We find that models exhibit high training accuracy in tasks such as $3/5$-people puzzles. The higher capacity GPT4o-mini nearly achieves interpolation ($\mathsf{Acc}(f;\mathsf{Tr})\approx 100\%$) using both Direct FT and CoT FT.
@@ 3.2 Fine-tuned Models
Figure 4: Train & test accuracy increases over the epochs. FTed LLMs can achieve interpolation ($\approx 100\%$ train accuracy) for easy tasks, e.g., 3/5-people puzzles. Llama3-8B struggles with CoT FT, likely due to its limited capacity.
4 LLMs Learn to Reason by Fine-tuning With Answers Only
§ 3 shows that both off-the-shelf and fine-tuned models exhibit memorization when solving K&K reasoning tasks. Does it mean that those models do not have reasoning capabilities at all? As we will show, it turns out that the models can do both, and interestingly the reasoning capability consistently improves as the memorization level increases when the models are fine-tuned on K&K puzzles.
We focus on analyzing Direct FT in this section and discuss CoT FT in § 5. For humans, solving K&K tasks without understanding the underlying logic is difficult. However, after observing the step-by-step reasoning steps, people can understand the procedure and solve the puzzles more easily. Similarly, compared to CoT FT, learning from only answers (Direct FT) without detailed reasoning steps is intuitively more challenging for LLMs, as the models need to come up with the reasoning procedures on their own. Therefore, the models might be more likely to rely on memorization in this case. Surprisingly, from Fig. 5, we did not observe Direct FTed GPT4o-mini models exhibiting consistently higher memorization score than CoT FTed ones. It turns out that models can learn to reason K&K puzzles well directly from observing only question-answer pairs, as we will show in § 4.1. To better understand what the model actually learns through Direct FT, we conduct a probing analysis on model internals in § 4.3 and an ablation study with incorrect answers fine-tuning in § 4.3.
4.1 Reasoning capabilities of Direct FT-ed model
Generalization performance increases with memorization level. As shown in Fig. 6 , the test accuracy ($y$-axis) of fine-tuned LLMs on the unseen test set continues to increase over the epochs, despite that the memorization score $\mathsf{LiMem}(f;\mathsf{Tr})$ on training samples also increases.
Figure 6: Test accuracy of fine-tuned Llama3-8B increases over epochs, despite the stronger memorization of training data as reflected by larger $\mathsf{LiMem}(f;\mathsf{Tr})$. See Fig. 18 for results on GPT4o-mini.
$\mathsf{LiMem}(f;\mathsf{Tst})$ on test samples is smaller than $\mathsf{LiMem}(f;\mathsf{Tr})$ on train samples in Fig. 5, particularly for more challenging (e.g., 5/8-people) tasks. Being able to relatively consistently solve the test puzzles after local perturbations, along with reasonable test accuracy in Fig. 5, suggests that models likely have learned to reason K&K puzzles to some extent and rely on reasoning when solving unseen test samples.

### v2

@@ Abstract
Large language models (LLMs) achieve good performance on challenging reasoning benchmarks, yet could also make basic reasoning mistakes. This contrasting behavior is puzzling when it comes to understanding the mechanisms behind LLMs’ reasoning capabilities. One hypothesis is that the increasingly high and nearly saturated performance on common reasoning benchmarks could be due to the memorization of similar problems. In this paper, we systematically investigate this hypothesis with a quantitative measurement of memorization in reasoning tasks, using a dynamically generated logical reasoning benchmark based on Knights and Knaves (K&K) puzzles. We find that LLMs could interpolate and memorize the training puzzles (achieving near-perfect accuracy) after fine-tuning, yet they struggle with slight variations of these puzzles. On the other hand, we show that while fine-tuning leads to heavy memorization, it also consistently improves generalization performance. Through in-depth analyses with perturbation tests, cross difficulty-level transferability, probing model internals, and fine-tuning with wrong answers, we establish that LLMs develop reasoning skills on K&K puzzles alongside memorization. Finally, our analysis based on a per-sample memorization score sheds light on how LLMs switch between reasoning and memorization when solving logical puzzles.
Machine Learning, ICML
@@ 1 Introduction
Modern Large Language Models (LLMs) show impressive reasoning capabilities that allow them to solve a wide range of challenging problems including commonsense reasoning and mathematical reasoning. In the meantime, LLMs also make mistakes on some of the most basic problems, such as comparing which number is bigger—13.11 or 13.8 (Lin, 2024), and counting the number of sisters that Alice’s brother has (Nezhurina et al., 2024). This contrast is puzzling when it comes to understanding how exactly LLMs solve reasoning tasks. This question is important both scientifically and practically: understanding how LLMs reason could shed light on their learning and generalization behaviors. It is also crucial for real-world applications where robust reasoning is required due to safety and trustworthiness concerns (Wang et al., 2023a; Wallace et al., 2024; Lee et al., 2024; Wei et al., 2024a).
@@
One hypothesis is that LLMs could be relying on memorization when solving those reasoning tasks, especially when measured by popular benchmarks that could be accidentally leaked into various massive internet-crawled pre-training datasets. Previous work (Tirumala et al., 2022; Carlini et al., 2023) show that LLMs could indeed memorize the training data, which may lead to potential privacy (Carlini et al., 2021) or copyright (Karamolegkou et al., 2023; Wei et al., 2024b) concerns. Additional evidences of potential memorization come from extensive studies on data contamination in LLMs (Magar & Schwartz, 2022; Balloccu et al., 2024; Shi et al., 2024; Xu et al., 2024; Oren et al., 2024). To mitigate the issue of benchmark saturation potentially due to memorization, some papers focus on designing dynamic benchmarks (Roberts et al., 2023; Zhu et al., 2024; Srivastava et al., 2024; Jain et al., 2024; Wu et al., 2024a) or alternative evaluation protocols (Zeng et al., 2023; Zhang et al., 2024; Xu et al., 2024; Srivastava et al., 2024).
Figure 1: Illustration of the definition of Local Inconsistency based Memorization Score, $\mathsf{LiMem}$. High level of memorization occurs when the model shows high accuracy in solving some problems but fails to consistently solve those problems under local perturbations that require similar underlying reasoning principles.
@@
In this paper, we take a direct approach to quantify the memorization behaviors of LLMs in reasoning tasks within a controlled setting. Specifically, we seek to understand: (i) whether LLMs rely on memorization to solve reasoning tasks, and (ii) whether memorization is only detrimental to learning to reason. Both questions are inspired by human behavior. For instance, when a student works hard on the preparation material for an exam, the preparation could help them get familiarized with the problems, and their ability to solve new problems could usually improve with enough exercises. However, without genuinely understanding the principles, they might fail when the same problem is slightly changed despite doing well on prepared problems. Our metric of memorization $\mathsf{LiMem}$, illustrated in Fig. 1, is based on this intuition: an LLM shows a high level of memorization when it solves reasoning problems with high accuracy but struggles to consistently solve those problems under local perturbations requiring similar mathematical principles (i.e., low consistency). We note that a similar perturbation (mostly at language-level) idea has been used in previous work, especially in detecting contamination (Golchin & Surdeanu, 2023; Yang et al., 2023; Xu et al., 2024). However, given our focus on understanding memorization in logical reasoning tasks, we further consider problem-level perturbation that slightly changes the mathematical structure of a puzzle, in addition to language-level perturbations. To facilitate our study, we propose a new logical reasoning benchmark that supports automatic problem-level perturbation. With this tool, we evaluate the reasoning power of 17 off-the-shelf LLMs. We then fine-tune Llama3-8B and GPT4o-mini to quantify their memorization in reasoning tasks, and reveal interesting interplay: while models indeed tend to memorize many training logical puzzles, they also develop reasoning capabilities during fine-tuning (even directly on question-answer pairs without reasoning steps), and the reasoning performance improves when memorizing more training puzzles.
In the following, we summarize our key contributions:
@@
  • To facilitate the measurement, we propose a new logical reasoning benchmark based on the Knights and Knaves (K&K, Smullyan, 1978; Johnson-Laird & Byrne, 1990) puzzles, that can generate new puzzles at different difficulty levels, locally perturb existing puzzles, and automatically synthesize detailed reasoning steps to solve a given puzzle (§ 2.2).
@@
  • We design various in-depth analyses (§ 4.1$\sim$ 4.3) to verify that LLMs developed improved reasoning capabilities (i.e., generalization) after fine-tuning even with only question-answer pairs, via local perturbation tests, cross difficulty-level transferability, fine-tuning with wrong answers, and model internal probing.
@@
2 Measuring Memorization in Reasoning
@@ 2.1 Memorization Metrics for Reasoning Tasks
Figure 2: K&K data generation framework employs abstract module and natural language module to generate question-answer pair and synthetic CoTs for each K&K puzzle, based on the problem specification: number of persons ($N$), tree width ($W$), and depth ($D$). Perturbers in these modules can alter the math structure and language description, respectively, and recompute the question-answer pair.
Figure 3: Test acc of off-the-shelf LLMs under 0-shot direct prompting drops with increasing puzzle complexity. For reference, OpenAI o1 model with test-time compute achieves an acc of 0.86 on 8-ppl task and 0.67 on 18-ppl task.
Memorization of LLMs has been studied in various contexts such as privacy (Carlini et al., 2023), copyright  (Carlini et al., 2021; Karamolegkou et al., 2023; Wei et al., 2024b; He et al., 2024), and solving knowledge intensive tasks (Hartmann et al., 2023). In this paper, we are specifically interested in measuring the level of memorization when solving reasoning tasks, by borrowing intuition from human behavior. For example, when preparing for an exam, a student may not be able to fully digest the underlying principles due to various reasons or constraints. But when (luckily) facing the same problem the student had prepared for, they would still be able to solve it. A key characteristic of this type of memorization is: (A) high accuracy on observed problems and (B) low accuracy when the problem is slightly changed. Based on this intuition, for a dataset $\mathcal{D}$ of reasoning puzzles, we combine the following two quantities to measure memorization:
@@
    For (A), we measure the accuracy of a target model $f$ on $\mathcal{D}$, denoted as $\mathsf{Acc}(f;\mathcal{D})$. We are especially interested in measuring on the set of observed puzzles, i.e., the training set, $\mathsf{Acc}(f;\mathsf{Tr})$. We say $f$ interpolates (Belkin et al., 2018; Muthukumar et al., 2020; Belkin, 2021; Bartlett et al., 2021) the training puzzles if $\mathsf{Acc}(f;\mathsf{Tr})\approx 100\%$.
@@
We combine the two factors to define a Local Inconsistency-based Memorization Score $\mathsf{LiMem}(f;\mathcal{D})\in[0,1]$:
$\mathsf{LiMem}(f;\mathcal{D})$ $=\mathsf{Acc}(f;\mathcal{D})\cdot(1-\mathsf{CR}(f;\mathcal{D}))$ (1)
=#$=\frac{\texttt{\#Correct}-\texttt{\#Consistently\_Correct}}{\texttt{\#Total}}.$
@@
When there is no ambiguity, we call it the memorization score. A larger score provides stronger evidence of memorization (i.e., a larger proportion of memorized examples in the given dataset). Specifically, a high $\mathsf{LiMem}(f;\mathsf{Tr})$ matches the characteristic behavior of human memorizing observed puzzles, and in this case we say $f$ memorized the training puzzles. Note that the $\mathsf{Acc}(f;\mathcal{D})$ factor is necessary, as there can be three types of behaviors: (i) solving by memorization, (ii) solving by reasoning, (iii) not solving (e.g., random guessing). A high $\mathsf{LiMem}(f;\mathcal{D})$ indicates (i), but a low $\mathsf{LiMem}(f;\mathcal{D})$ would only indicate (ii) if we separately check that $\mathsf{Acc}(f;\mathcal{D})$ is high.
@@
To effectively measure the memorization score $\mathsf{LiMem}(f;\mathcal{D})$, we need a principled way to (1) locally perturb the puzzle while maintaining its difficulty level; (2) compute the new correct answer after perturbation. Towards this goal, we design and implement a functional dataset based on the Knights and Knaves puzzles (Smullyan, 1978; Johnson-Laird & Byrne, 1990).
@@
Figure 4: Train & test accuracy increases over the epochs. FTed LLMs can achieve interpolation ($\approx 100\%$ train accuracy) for easy tasks, e.g., 3/5-ppl puzzles. Llama3-8B struggles with CoT FT on K&K tasks, likely due to limited model capacity.
Figure 5: Fine-tuned LLMs generally exhibit both higher clean accuracy (x-axis) & inconsistency ratio under perturbations (y-axis) on the train set than test set, resulting in a higher memorization score (color spectrum). LLMs show stronger memorization under math-level perturbations (statement/leaf) than language level. We separately report memorization score in Fig. 23 and consistency ratio in Fig. 24, and results under combined math & language-level perturbations in Fig. 25.
2.2 Knights and Knaves Reasoning Benchmark
Knights and Knaves (K&K) is a type of logical puzzle, where the goal is to infer each character $i$’s truthfulness $B_{i}$ (Boolean value) by judging the logical consistency of the statements $S_{i}$ they made. Fig. 3 shows an example.
The principle underlying K&K is the Boolean satisfiability problem (SAT) (Boolean satisfiability problem, ). SAT was the first problem proven to be NP-complete and many well-known problems can be translated into SAT, such as hardware and software verification and theorem proving (SAT solver, 2024). Hence, the performance of a model on SAT (i.e., K&K puzzles) can be important indicative of its reasoning capabilities. Specifically, consider a K&K puzzle involving $N$ people, and a possible Boolean value assignments to $\{B_{i}\}_{i=1}^{N}$, where $B_{i}$ indicates whether the $i$th person is telling the truth, i.e., their statement $S_{i}$ is true. Therefore, a valid solution to a K&K puzzle is an assignment such that the following formula is true: $(B_{1}\Leftrightarrow S_{1})\wedge(B_{2}\Leftrightarrow S_{2})\wedge\cdots\wedge(B_{N}\Leftrightarrow S_{N})$.
Based on the K&K puzzle, we design a dynamic benchmark that supports generating new puzzles and local perturbations. Our benchmark has 2 modules (See Fig. 3 for an overview):
The Abstract Module contains 4 components that generate and manipulate K&K puzzles in an abstract form: The Generator that produce random puzzles; the Solver that find valid solutions algorithmically; the Reasoner that generate human-like reasoning steps (chain-of-thoughts, CoT); and the Perturber that maps a given puzzle to a local perturbation. Each puzzle involves $N$ people, each making a statement forming a logical tree with max width/depth of $W$/$D$, using the logical operations and, or, not, implication, and equivalence. The Perturber is the most important part that is not usually supported in existing benchmark. It generates perturbation by replacing a statement or a leaf node in a statement with a newly sampled one, and ensures the perturbation has a different solution. This support is crucial for making our memorization measurements.
The Natural Language Module converts the puzzles and the generated CoTs into natural language. It uses random names and templates to diversify the generated puzzles, and also supports language level perturbation to a given puzzle.
See § C for more details. We generate and release a core dataset of $1000$/$100$/$50$ train/test/validation puzzles for each $2\leq N\leq 8$ people [^1] We only include $200$ training examples for 2-ppl puzzles due to the limited problem space. Note the problem space is huge as $N$ increases. For example, for 8-ppl puzzles ($D,W=2,2$), there are ~$10^{24}$ unique problems, and ~$30\%$ of them has a unique solution based on empirical estimation from 100k random generations.. By default we use a maximum tree width of $W=2$ and depth $D=2$. For each puzzle, we also generate six perturbed variants (2 problem level and 4 language level): {perturbed statement, perturbed leaf node, random role-pair name, uncommon person name, reordered statement, flipped role}. We release the code for generation additional puzzles and perturbations if needed.
K&K is challenging for off-the-shelf models. We use 0-shot direct prompting with task-specific instructions for open-ended question-answering (details in § D.2). [^2] Note that even under direct prompting, capable LLMs can generate Chain of Thought (Wei et al., 2022, CoT). Our evaluation mainly considers the 00-shot setting to avoid biases from in-context examples (Zhao et al., 2021), but we provide results for CoT prompting, 1-shot prompting & self-consistency prompting in § E. Accuracy is determined by keyword matching and requires correctly identifying all characters in the conclusion. We evaluate 17 leading models known for strong reasoning performance. Fig. 3 shows that K&K puzzles are highly challenging—even for the simplest 2-ppl puzzles, the best models (except o1 model) achieve at most 70% accuracy, which drops to just 11% for 8-ppl puzzles. In § E.1, we show that various prompting techniques like CoT/1-shot/self-consistency (Wang et al., 2023b) cannot fundamentally improve performance on challenging K&K tasks.